OpenStudy (anonymous):

Use the Intermediate Value Theorem to show that the equation x=cos x has at least one solution in the interval [0, pi/2].

jimthompson5910 (jim_thompson5910):

hint: turn x = cos(x) into cos(x) - x = 0

OpenStudy (anonymous):

oh, lemme do that really quick.

jimthompson5910 (jim_thompson5910):

let f(x) = cos(x) - x

jimthompson5910 (jim_thompson5910):

then evaluate f(0) and f(pi/2)

OpenStudy (anonymous):

Thank you, i understand now.

jimthompson5910 (jim_thompson5910):

yw